OpenStudy (anonymous):

Factor and then solve to find the roots of the equations x^2-6x = 0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2-6x+%3D+0

OpenStudy (asadkarim7):

X^2= 6x x=+/- 6

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

\[x(x-6)=0\] \[x=0\] or \[x-6=0\] \[x=6\]

OpenStudy (anonymous):

solutions are \[\{0,6\}\]